### v1

# Self-Tuning Spectral Clustering for Speaker Diarization

Abstract
Spectral clustering has proven effective in grouping speech representations for speaker diarization tasks, although post-processing the affinity matrix remains difficult due to the need for careful tuning before constructing the Laplacian. In this study, we present a novel pruning algorithm to create a sparse affinity matrix called spectral clustering on p-neighborhood retained affinity matrix (SC-pNA). Our method improves on node-specific fixed neighbor selection by allowing a variable number of neighbors, eliminating the need for external tuning data as the pruning parameters are derived directly from the affinity matrix. SC-pNA does so by identifying two clusters in every row of the initial affinity matrix, and retains only the top $p\%$ similarity scores from the cluster containing larger similarities. Spectral clustering is performed subsequently, with the number of clusters determined as the maximum eigengap. Experimental results on the challenging DIHARD-III dataset highlight the superiority of SC-pNA, which is also computationally more efficient than existing auto-tuning approaches.
Index Terms:
speaker diarization, spectral clustering, matrix sparsification, eigengap, DIHARD-III
I Introduction
The availability of precise annotations of audio recordings based on speaker information can significantly enhance the field of audio analytics involving multi-speaker conversations. This has applications across various domains, from everyday online meetings to advanced speech and audio forensics. This task is formally known as speaker diarization (SD), where the goal is to assign speaker labels to segments of speech based on speaker identity [1]. Traditionally, SD addresses the problem of “who spoke when” in a given audio recording, particularly in multi-speaker environments. SD pipeline conventionally consists of several interconnected components, including speech enhancement (SE), speech activity detection (SAD), segmentation, speaker embedding extraction, clustering, and re-segmentation.
Modern SD architectures based on deep learning [2] have demonstrated remarkable performance, surpassing classical methods on widely used benchmark datasets such as CALLHOME [3], AMI [4], and VoxSRC [5]. However, SD systems still face practical challenges when applied to more realistic conversational speech datasets, such as subsets of the DIHARD dataset [6]. Factors like background noise, overlapping speech, a large number of speakers, imbalanced speech contributions across speakers, intra-speaker variability, and the recording environment’s lack of domain adaptation make the task even more complex. While issues such as environmental noise, SAD, speech overlap, and intra-speaker variability are extensively studied, the clustering problem does not receive the necessary attention despite playing a crucial role in the overall diarization process.
Clustering is a classical problem in unsupervised machine learning with a substantial body of literature [7, 8]. However, in the context of speaker diarization (SD), clustering algorithms are largely limited to basic techniques such as k-means, agglomerative hierarchical clustering, mean-shift [9], and spectral clustering (SC) [10]. Among these, SC is the most widely used, particularly in modern SD systems based on deep speaker embeddings [11]. This preference is primarily due to its simplicity, with fewer parameters, and its strong mathematical foundations [12, 13].
The spectral clustering approach used for the SD task involves specific parameters during the adjustment of the affinity matrix. Since the nodes represented by speaker embeddings capture speaker characteristics, it is crucial to minimize the impact of different speaker similarities and to retain or enhance same-speaker similarities [11]. In practice, a sparse graph is formed by removing less similar nodes and setting their values to zero. This process requires a pruning parameter typically tuned using externally labeled speech data. However, audio from different domains may introduce domain mismatches, significantly degrading SD performance. Recently, an auto-tuning approach for SC, named as auto-tuning spectral clustering (ASC), has been proposed in the context of SD, eliminating the need for tuning on external data [11]. The tuning parameters are instead derived directly from the audio recording being evaluated, using a brute-force search across different parameter values and optimizing a proxy evaluation metric. While this technique is promising, it has several limitations. First, it requires repeated spectral clustering computations involving eigenvalue decomposition, making it computationally expensive for longer recordings. Second, the proxy evaluation metric has limitations and does not always accurately reflect actual diarization performance [11]. Third, it uses a fixed number of neighbors for each node, which is problematic, especially for audio recordings with an unbalanced distribution of speech. Additionally, the graph may become disconnected in specific situations, leading to the failure of the spectral clustering algorithm or suboptimal results [13, 14].
This work introduces a novel self-tuning method called Spectral clustering on p-neighborhood retained affinity matrix (SC-pNA) for speaker diarization, addressing the limitations of previous state-of-the-art (SOTA) adaptive spectral clustering approaches. Inspired by the score distribution characteristics reported in speaker recognition literature [15, 16], we propose a score-based, node-specific pruning strategy. In this method, the pruning threshold is dynamically adjusted for each node, allowing the retention of nodes with stronger connections while eliminating weakly connected, unreliable nodes. Like ASC, our approach does not require external data for threshold adjustment. However, it advances further by eliminating the need for proxy evaluation metrics. Additionally, it requires only a single eigenvalue decomposition of the Laplacian matrix, which is computed once from the refined adjacency matrix.
II Spectral clustering for speaker diarization
II-A Conventional spectral clustering (CSC)
Spectral clustering  [13] is a technique used to group data points into clusters based on the eigenvalues and eigenvectors of a similarity matrix constructed from the data. It leverages the properties of the graph Laplacian to partition the data in a way that often captures complex cluster structures better than traditional methods like k-means. It has considerable applications in the area of image segmentation [14]. With the emergence of speech embeddings [17, 18, 19], SC is widely adopted for SD [10, 20, 21, 2]. The SC framework for SD is summarized as follows:
Given a finite set $\mathbf{X}=$$x_{1},x_{2},...,x_{N}$} of $N$ embeddings extracted from their corresponding $N$ speech segments. Our objective is to cluster them into $k$ speaker classes. First, an affinity matrix $\mathbf{M}\in\mathbb{R}^{N\times N}$ is computed. Each element in $\mathbf{M}$ represent the cosine similarity between speech embeddings. Unreliable similarity scores are pruned in each row of $\mathbf{M}$. A pruning parameter $\alpha$ determines the fixed $\left\lfloor N(1-\alpha)\right\rfloor$ entries to be made zero. The resulting matrix $\mathbf{M}_{\alpha}$ is symmetrized. Further, an unnormalized Laplacian matrix is computed as $\mathbf{L}=\mathbf{D}-\mathbf{M}_{\alpha}$, where $\mathbf{D}$ is a degree matrix. The standard $k$-means algorithm is used to cluster the first $k$ eigenvectors corresponding to higher eigenvalues obtained from the eigendecomposition of $\mathbf{L}$. The optimal pruning parameter $\alpha$ is estimated by tuning it on the development set. It is done by varying $\alpha$ linearly from 0 to 1 with a typical step size of 0.01. The value of $\alpha$ that exhibits the minimum diarization error rate (DER) is utilized for the evaluation set. But this setting is not the best choice, as in scenarios such as where the development and evaluation data can come from situations that are substantially different [22].
II-B Auto-tuning spectral clustering (ASC)
To mitigate this data-dependency, another variant of SC known as auto-tuning spectral clustering (ASC) was proposed in [11]. This approach shares a fundamental similarity with CSC, as both require repeated computation of spectral clustering (SC) at varying pruning rates. However, rather than relying on actual DER measured with the ground truth from a labeled development set, this method estimates a proxy DER directly from the unlabelled data being diarized. The proxy DER is calculated by first determining the maximum eigengap, normalized by the largest eigenvalue, and then computing the ratio of the pruning factor to the maximum eigengap.
This technique is preferred over the CSC as it does not require any ground truth but it also has some limitations. It is computationally expensive, as it requires eigen-decomposition of the Laplacian matrix $\mathbf{L}$ as many times as the number of embeddings $N$. So, it is not a realistic choice in scenarios where the recordings are arbitrarily longer. Also, it works on the concept of a proxy-DER, which may always not be a good approximation of the original DER. Also, it prunes a fixed number of elements in each row of the affinity matrix, which is also not an ideal choice in situations when the distribution of spoken duration of speakers is not uniform across a recording.
II-C Spectral clustering on $p$-neighborhood retained affinity matrix (SC-pNA)
Background: Our proposed method of SC-pNA is focused on creating a sparse affinity matrix that can aid the spectral clustering method to better delineate different speaker embeddings, thereby improving speaker diarization. To create this sparse affinity matrix, each row is pruned independently by removing low similarity scores. Each row of the original affinity matrix should ideally have high similarities to the embeddings of the same speakers, while having low similarities to the embeddings of all other speakers.
As the number of embeddings belonging to the same speaker will vary across a recording, therefore the number of similarity scores retained in each row should be decided adaptively. In general, the similarity scores in each row of an affinity matrix can be described in terms of two distributions. The first within-cluster distribution $C_{w}$ corresponds to the high similarity scores between same speaker embeddings, whereas the second between-cluster distribution $C_{b}$ describes the low similarity scores between different speaker embeddings. This approach of describing similarities in terms of two distributions has been previously explored in the context of speaker recognition [15, 23], where $C_{w}$ and $C_{b}$ are assumed to follow the Gaussian distribution, i.e., $C_{w}\sim\mathcal{N}(\mu_{w},\sigma_{w})$, and $C_{b}\sim\mathcal{N}(\mu_{b},\sigma_{b})$. As we wish to accept $C_{w}$ and reject $C_{b}$, the equal error rate (EER) provides a location where the false rejection rate of $C_{w}$ equals the false acceptance rate of $C_{b}$. The analytical EER is defined as,
$\text{EER}=\frac{1}{2}-\frac{1}{2}\text{erf}\left(\frac{\text{F-ratio}}{\sqrt{2}}\right),$ (1)
where erf(.) is the error function for Gaussian distributions, and the F-ratio is $(\mu^{w}-\mu^{b})/(\sigma^{w}+\sigma^{b})$. The optimal threshold corresponding to the EER, shown in Figure 1, is given by,
$\Delta=\frac{\mu^{w}\sigma^{b}+\mu^{b}\sigma^{w}}{\sigma^{w}+\sigma^{b}}.$ (2)
The theoretical foundations of the threshold $\Delta$ make it an appealing choice to delineate distributions $C_{w}$ and $C_{b}$ in each row of the affinity matrix, and thereafter all entries corresponding to $C_{b}$ can be set to zero to create the sparse affinity matrix. We refer to this method as the EER-$\Delta$ approach: on an initial affinity matrix $A$, $\Delta_{i}$ is first computed over each row $A_{i}$, for $i=1,...,n$. A sparse affinity matrix $B$ is then created by using $\Delta_{i}$ as a threshold in each row $B_{i}$, where we set $B_{ij}=0$ if $A_{ij}<\Delta_{i}$, otherwise $B_{ij}=A_{ij}$. Finally, spectral clustering is conducted on $B$, using the maximum eigengap to determine the number of clusters. The EER-$\Delta$ approach has a lower computation cost compared to the CSC and ASC as it does not require tuning. An additional advantage of EER-$\Delta$ is that the number of similarity scores pruned in each row can vary, depending on the sizes of the identified $C_{w}$ and $C_{b}$.
[…]
Figure 1: The EER-$\Delta$ approach accepts all similarity scores in cluster $C_{w}$ up to the EER optimal $\Delta$ threshold, whereas the proposed SC-pNA only retains the top $p\%$ similarity scores in cluster $C_{w}$, and prunes all lower similarities.
SC-pNA Methodology: In SC-pNA, we consider an even more aggressive thresholding approach, to create a sparser affinity matrix. As the $\Delta_{i}$ thresholds of EER-$\Delta$ retain all similarity scores for $C_{w}$ in the $i$-th row, we can interpret this approach as restricting the neighborhood of the $i$-th speaker segment $x_{i}$ only to the $C_{w}$ cluster. Let $N(x_{i})$ be the set of speaker segments in the neighborhood of $x_{i}$, and let $C_{w}^{(i)}$ be the cluster identified for $x_{i}$. Then each $\Delta_{i}$ threshold of EER-$\Delta$ establishes,
$N(x_{i})=C_{w}^{(i)}\text{ for }i=1,2,...,n,$ (3)
throughout the affinity matrix.
In the proposed SC-pNA, a more aggressive pruning is pursued, where only the top $p\%$ similarities from the identified $C_{w}^{(i)}\backslash\{x_{i}\}$ are retained in every row. This is shown in Figure 1, where the proposed pruning approach can be interpreted as considering higher false rejection ratio (FAR), with the thresholds empirically set at (1-p)FAR. The constant high self-similarity scores $s(x_{i},x_{i})=1$ are excluded from the process to identify the clusters $C_{w}^{(i)}$ and $C_{b}^{(i)}$, as they can behave as outliers when all other similarities $s(x_{i},x_{j}),\forall j\neq i$, are much smaller than $s(x_{i},x_{i})$. This exclusion is interpreted as the removal of all self-loops, thus leading to an adjacency matrix where connectivity is maintained only through neighborhood connections. Thus retaining only top $p\%$ similarities creates the neighborhoods for all $x_{i}$,
$N_{p}(x_{i})=S^{(i)}\subseteq C_{w}^{(i)}\backslash\{x_{i}\}.$ (4)
Comparing equations (3) and (4) lets us determine that $|N_{p}(x_{i})|\leq|N(x_{i})|$, and thus smaller neighborhoods are formed, leading to more sparse affinity matrices. However, these smaller neighborhoods should be capable of identifying the correct speaker clusters. If a recording contains $k$ speakers, the embeddings should be clustered correctly into $k$ clusters $C_{1},\ldots,C_{k}$, where each cluster $C_{j}=\{x_{j1},x_{j2},\ldots,x_{j|C_{j}|}\}$ contains $|C_{j}|$ embeddings. Thus $p$ should be chosen so that $|N_{p}(x_{i})|$ is small, yet cluster $C_{j}$ can be recovered by examining the neighborhoods of $\{x_{j1},x_{j2},\ldots,x_{j|C_{j}|}\}$, i.e.,
$N_{p}(x_{j1})\cup N_{p}(x_{j2})\cup\ldots\cup N_{p}(x_{j|C_{j}|})=C_{j}.$ (5)
In the EER-$\Delta$ approach, $N(x_{j1})\cup\ldots\cup N(x_{j|C_{j}|})=C_{j}$ holds trivially. For real world data, that can be noisy and have overlapped clusters, the challenge for both SC-pNA and the EER-$\Delta$ approach lies in utilising the available neighborhood information to identify the set of embeddings $\{x_{j1},x_{j2},\ldots,x_{j|C_{j}|}\}$ that recovers each cluster $C_{j}$ accurately.
The following are the steps of the proposed SC-pNA approach:
  1. 1.
    Create the adjacency matrix without self-loops: The initial adjacency matrix $A$ is formed using cosine similarity,
    $A_{ij}=\frac{x_{i}^{T}x_{j}}{||x_{i}||\,||x_{j}||},\ \forall i,j=1,2,\ldots,n.$ (6)
    The diagonal entries of $A$ are set to zero i.e., $A_{ii}=0,\ i=1,2,\ldots,n$.
  2. 2.
    Identify $C_{w}^{(i)}\ \text{and}\ C_{b}^{(i)}$ in each row: In each row $A_{i}$, two clusters $C_{1}^{(i)},C_{2}^{(i)}$ are identified using $k$-Means clustering with $k=2$. From these two clusters, $C_{w}^{(i)}$ is identified as the one with the larger cluster center,
    $C_{w}^{(i)}=\text{argmax}\{\text{center}(C_{1}^{(i)}),\text{center}(C_{2}^{(i)})\}.$ (7)
    $C_{b}^{(i)}$ is identified as the other cluster i.e., $C_{b}^{(i)}=\text{argmin}\{\text{center}(C_{1}^{(i)}),\text{center}(C_{2}^{(i)})\}$.
  3. 3.
    Retain top $p\%$ similarity scores: A matrix $P$ is formed which retains in each row the top $p\%$ similarity scores from the cluster $C_{w}^{(i)}$, and the rest are set to zero, i.e.,
    $P_{ij}=\begin{cases}0&\text{, if }A_{ij}\text{ not in the top }p\%\text{ of }C_{w}^{(i)}\\ A_{ij}&\text{, otherwise}\end{cases}.$ (8)
  4. 4.
    Symmetrize and form the Laplacian: A symmetric matrix $W$ is formed from $P$,
    $W=\frac{1}{2}(P+P^{T}).$ (9)
    A diagonal matrix $D$ is created with the sum of the rows of $W$ as its entries: $D_{ii}=\sum_{j=1}^{n}|w_{ij}|$. The Laplacian $L$ is then created,
    $L=D-W.$ (10)
  5. 5.
    Compute eigengap and estimate $\hat{k}$ : For each recording, a user can specify a reasonable assumption for the maximum number of speakers $k_{\max}$. We find the smallest $M$ eigenvalues $\lambda_{1},\ldots,\lambda_{M}$ of $L$, where $M=\min\{k_{\max},n\}$. The $M-1$ length eigengap vector is constructed as,
    $\mathbf{e_{gap}}=[\lambda_{2}-\lambda_{1},\lambda_{3}-\lambda_{2},\dots,\lambda_{M}-\lambda_{M-1}].$ (11)
    The number of speakers $\hat{k}$ is then estimated as,
    $\hat{k}=\text{argmax}\ \mathbf{e_{gap}}.$ (12)
  6. 6.
    Cluster the spectral embeddings: Form matrix $V\in\mathbb{R}^{n\times\hat{k}}$, containing in its columns the $\hat{k}$ eigenvectors corresponding to the smallest $\hat{k}$ eigenvalues of $L$. Apply, $k$-Means clustering on the rows of $V$ to obtain cluster memberships $\hat{y}\in[1,2,\ldots,\hat{k}]^{n}$.
The algorithm for the proposed SC-pNA is outlined in Algorithm 1. As SC-pNA does not require any tuning, and only performs a single eigenvalue decomposition of the Laplacian matrix, it incurs significantly lower computation costs compared to the SOTA ASC approach.
Algorithm 1 SC-pNA
  Input: Set of embeddings $\mathbf{X}\in\mathbb{R}^{n\times d},\ p$.
  Output: Cluster memberships $\hat{y}\in[1,2,\ldots,\hat{k}]^{n}$.
  1. Compute $A$ using equation (6).
  2. for $i=1$ to $n$
  3.  Set $A_{ii}:=0$.
  4.  Form clusters $C_{1}^{(i)}$, $C_{2}^{(i)}$ using $k$-Means with $k:=2$.
  5.  Identify $C_{w}^{(i)}$ using equation (7).
  6. Form $P$ using equation (8).
  7. Form $W$ using equation (9).
  8. Form $L$ using equation (10).
  9. Compute the $\mathbf{e_{gap}}$ using equation (11).
  10. Estimate $\hat{k}$ using equation (12).
  11. Obtain $\hat{y}$ using $k$-Means on the spectral embeddings in $V$, corresponding to the $\hat{k}$ smallest eigenvalues of $L$.
III Experimental Setup
We empirically compare the unsupervised speaker diarization performances of the proposed SC-pNA (with $p=20\%$ retention), with the EER-$\Delta$ approach, and the SOTA method of ASC. As a baseline, the semi-supervised speaker diarization performance of CSC is also measured. We also empirically validate the robustness of SC-pNA by observing its performance over a range of retention percentages $p$ from $10$ to $50$, with increments of $5$.
III-A Dataset
For our experiments, we utilized the speech corpora from the third DIHARD speech diarization challenge (DIHARD-III) [6]. It comprises data from eleven diverse domains characterized by a variety in the number of speakers, conversation types and speech qualities. For this study, we focused on seven domains: broadcast interview, court, cts, maptask, meeting, socio lab, and webvideo. These domains are representative of everyday conversational speech scenarios between adults. We excluded four domains because they either contained single-speaker data (audiobooks) or regions with personally identifiable information (PII) (clinical, restaurant, and socio field data). For each domain there exists two splits of development (dev) and evaluation (eval). Both splits are used to evaluate the SD methods, as the methods operate in an unsupervised manner. This results in a total of fourteen data splits over the seven data domains.
III-B SD system
We conducted our experiments using the SpeechBrain toolkit [24]. Our implementation is based on a modified version of the AMI recipe provided in the toolkit  [^1] https://github.com/speechbrain/speechbrain/tree/develop/recipes/AMI, which employs a pre-trained ECAPA-TDNN model as the speaker embedding extractor [25]. The ECAPA-TDNN model was trained on the VoxCeleb dataset [26, 27]. The extracted speaker embeddings are 192-dimensional, derived from the penultimate layer of the ECAPA-TDNN architecture. To ensure robust performance and eliminate the variability introduced by SAD, we utilized ground truth annotations. Each segment is set to a duration of 3.0 seconds, with an overlap of 1.5 seconds between consecutive segments.
III-C Evaluation metric
We use standard DER mertric for evaluating the performance of the diarization system [28]. DER is comprised of three key errors: missed speech, false alarm of speech, and speaker error. DER is quantified as the ratio of the combined duration of these three errors to the total duration.
IV Results & Discussion
IV-A Comparison with SOTA
The unsupervised SD performances of the proposed SC-pNA (with $p=20\%$ retention) and the EER-$\Delta$ approach are compared with the SOTA method of ASC. The resulting DERs achieved over the seven domains of DIHARD-III are shown in Table I. We observe that SC-pNA achieves the lowest DER among the unsupervised methods over nine of the fourteen data splits. ASC achieved the lowest DER among the remaining four data splits, whereas EER-$\Delta$ in general obtained higher DERs across the data splits. With the exception of the court domain, SC-pNA has in general obtained better DER compared to ASC, and is competitive on the two splits where ASC has lower DER. Thus we can recommend SC-pNA for the task of unsupervised SD.
TABLE I: Comparison of DER (lower is better) achieved by the unsupervised SD methods of ASC, EER-$\Delta$, and SC-pNA, over seven domains of DIHARD-III. The CSC provides a semi-supervised SD baseline.
Domain Split CSC ASC EER-$\Delta$ SC-pNA
broadcast dev 2.03 4.09 2.41 2.98
interview eval 3.58 3.67 6.82 4.77
court dev 1.82 2.73 16.68 6.04
eval 2.09 2.73 17.51 7.15
cts dev 8.28 9.73 21.40 8.22
eval 6.58 7.3 12.66 6.63
maptask dev 2.19 4.8 10.05 2.73
eval 1.78 7.47 5.25 0.92
meeting dev 16.11 18.6 26.65 16.79
eval 16.79 19.60 40.84 21.26
socio lab dev 3.38 4.6 8.33 3.08
eval 1.99 4.56 8.66 1.97
webvideo dev 33.94 41.57 35.68 30.68
eval 36.29 37.60 36.52 33.06
Comparing the unsupervised SD performance of SC-pNA with the semi-supervised SD performance of CSC, we observe that with the exception of the court domain, SC-pNA in fact obtains DERs that are close to the DERs of CSC. For six of the data splits: cts (dev), maptask (eval), socio lab (dev and eval), and webvideo (dev and eval), the unsupervised SC-pNA even outperforms the semi-supervised CSC. Thus in general, the SD performance of SC-pNA is observed exceed the performance of the SOTA ASC, and is also competitive with respect to the semi-supervised method of CSC.
IV-B Variations in DERs for different $p\%$ in SC-pNA
The previous study involved SC-pNA with retention percentage $p$ as $20\%$. Here we study the variation in the resulting DER for values of $p$ in the range of $10\%$ to $50\%$, with an increment of $5\%$. For four data splits of maptask (eval), meeting (eval), socio lab(eval), and webvideo (eval), in Figure 2 the DERs obtained by SC-pNA are shown, and are compared with those obtained by CSC and ASC. From this figure, we observe that for maptask (eval) and webvideo (eval), the DERs obtained across all retention percentages are better than not only those obtained by ASC, by also those obtained by the semi-supervised CSC. For meeting (eval) the DER is generally higher than ASC or CSC, and we observe that at the retention percentages of $25\%$ and $20\%$ the DERs are the closest to ASC. For socio lab (eval) the DERs are also lower than ASC, and quite comparable to the baseline CSC. Thus we conclude that the proposed SC-pNA in general shows robust SD performance across different retention percentages, and the selected $20\%$ retention can be recommended as it generally achieves low DERs across various data splits.
Figure 2: For four data splits, the DERs (lower is better) obtained by SC-pNA over different retention percentages are compared with the DERs obtained by CSC and ASC.
V Conclusion
We proposed an SD approach called SC-pNA, which creates a sparse affinity matrix following an approach motivated by the theory on EER for Gaussian distributions. SC-pNA identifies two clusters in each row of an initial affinity matrix, and retains the top $p\%$ similarity scores from the cluster with larger similarities. This approach enables SC-pNA to prune a variable number of similarity scores. The resulting sparse affinity matrix is symmetrized, from which the Laplacian is formed. The speakers are then identified by clustering the spectral embeddings, based on the number of speakers that is automatically selected from the eigengap. The overall method of SC-pNA has significantly lower computation cost compared to the SOTA ASC approach, while our empirical results show that SC-pNA with a retention percentage of $20\%$ outperforms the ASC over the DIHARD-III dataset, making SC-pNA a method to be recommended for SD. Empirical results also notably showed that SC-pNA was competitive against the semi-supervised CSC, and showed robust performances across different retention percentages.
Acknowledgment
The primary author expresses sincere gratitude to the Linguistic Data Consortium (LDC) for the LDC Data Scholarship, which enabled access to the DIHARD-III dataset.